OpenStudy (anonymous):

Can anyone help me with this? I need to find the domain and range for this equation g(x)=(x^2 - 2x + 5)^(1/2)

Parth (parthkohli):

What is the domain for the square root function?

OpenStudy (anonymous):

positive and negetive infinity

Parth (parthkohli):

Eh? I think it should be more like \((+\infty , 0]\)

Parth (parthkohli):

So the thing inside the square root should be non-negative right?

OpenStudy (anonymous):

yeah

Parth (parthkohli):

\[x^2 - 2x + 5 \ge 0\]The thing you get for \(x\) is the domain.

OpenStudy (anonymous):

I got that so far but I'm having problem with what it can't be like x neq 3 or .6 etc.

Parth (parthkohli):

Hmm...

Parth (parthkohli):

\[x(x - 2) \ge -5\]true for all reals.

Parth (parthkohli):

so the domain is all reals.

OpenStudy (anonymous):

that's it?